# Adaptive Informed Deep Neural Networks for Power Flow Analysis

### v2

@@ Abstract
This study introduces PINN4PF, an end-to-end deep learning architecture for power flow (PF) analysis that effectively captures the nonlinear dynamics of large-scale modern power systems. The proposed neural network (NN) architecture consists of two important advancements in the training pipeline: (A) a double-head feed-forward NN that aligns with PF analysis, including an activation function that adjusts to the net active and reactive power injections patterns, and (B) a physics-based loss function that partially incorporates power system topology information. The effectiveness of the proposed architecture is illustrated through 4-bus, 15-bus, 290-bus, and 2224-bus test systems and is evaluated against two baselines: a linear regression model (LR) and a black-box NN (MLP). The comparison is based on (i) generalization ability, (ii) robustness, (iii) impact of training dataset size on generalization ability, (iv) accuracy in approximating derived PF quantities (specifically line current, line active power, and line reactive power), and (v) scalability. Results demonstrate that PINN4PF outperforms both baselines across all test systems by up to two orders of magnitude not only in terms of direct criteria, e.g., generalization ability, but also in terms of approximating derived physical quantities.
Index Terms:
@@ I Introduction
PF analysis can be performed by solving nonlinear and non-convex algebraic equations derived from the nodal balance of the net active and reactive power injections at each bus in power systems [3]. Exact analytical solutions for these equations, which also involve impedance parameters, load characteristics, and generator conditions of power systems, are impossible. Therefore, iterative numerical methods, such as the Gauss-Seidel and Newton-Raphson (NR) methods, are conventionally employed to converge to a solution that satisfies the PF equations within specified accuracy limits. Eventually, the solutions yield voltage phasors across the entire power system and, hence, provide a comprehensive understanding of the operational state of the power system [4, 5, 6].
@@
Several studies in the literature have investigated the impact of various modifications on the efficacy of deep learning-based PF analysis (e.g., [20, 21]). These modifications are categorized into three main stages: (i) pre-training (e.g., [22]); (ii) training (e.g., [23]); and (iii) post-training (e.g., [24]), as outlined in Table I. A majority of these studies have primarily focused on the training stage. The table also highlights instances where topology information, such as line physical properties (e.g., [25]) and graph topology (e.g., [26]) has been integrated into the training stage. The literature review also indicates the utilization of different prior knowledge for deep learning-based PF analysis and optimal power flow (OPF), including equality and inequality constraints (e.g., [27]) and PF equations (e.g., [28]). Among the three presented stages, this work focuses on the training stage.
TABLE I: List of the literature on deep learning-based power flow (PF) and optimal power flow (OPF). The scope of the studies is grouped into three stages: (i) pre-training; (ii) training; (iii) post-training.
@@
$p_{i}$=nodal active power, $q_{i}$=nodal reactive power, $s_{ij}$=line apparent power, $|v|$=voltage magnitude, $\delta$=voltage angle, $\mathcal{G}$=graph topology, $l_{ij}$=line physical properties, $A$=Adjacency matrix, $p^{g}$=generator active power, $q^{g}$=generator reactive power, $p^{d}$=load active power, $q^{d}$=load reactive power.
Measurement data has been used in [22] to enhance deep learning-based OPF. While this modification has demonstrated improvements in accuracy, its applicability is confined to the pre-training stage. Physical properties have been used on a few occasions in the past to enhance deep learning-based PF analysis (e.g., [25, 38]), where the focus was mainly on the training stage. However, the high dependency on the physical properties of power systems affects the efficiency of this modification for deep learning-based PF analysis, certainly because physical properties are not always reliable due to different reasons, including aging and environmental conditions. In addition, integrating all the physical properties makes training processes computationally very expensive. The completion idea and filtering technique have also been employed in a few studies (e.g., [37]), where the focus is mainly on the post-training stage. Yet, similar to pre-training stage modifications, this approach has limitations as it is detached from the training stage. Among various methods, modified loss functions have been widely used for deep learning-based PF analysis (e.g., [29]) and OPF (e.g., [34]). Physics-based loss functions have shown significant promise in enhancing the performance and reliability of NNs for PF and OPF.
This paper proposes an end-to-end deep learning architecture, hereafter called PINN4PF, to develop scalable and physics-informed NNs for PF analysis. The focus is on the training stage and extends the existing knowledge of physics-informed NNs for PF analysis, e.g., [23, 29]. The main contributions are:
@@
    developing a double-head feed-forward NN that aligns with PF analysis and includes an activation function that adjusts to the net active and reactive power injections patterns, i.e., input features, and
@@
    developing a physics-based loss function that partially incorporates power system topology information.
The proposed PINN4PF architecture is designed to predict complex bus voltages in power systems by learning the mapping from the net active ($p$) and reactive ($q$) power injections to voltage states. It employs a double-head feed-forward NN structure, where one head predicts the real component ($\mu$) and the other predicts the imaginary component ($\omega$) of the complex voltages. Each head utilizes an adaptive activation function, a modified ReLU, whose slope coefficient is a trainable parameter. This design allows the activation functions to adapt dynamically during training via backpropagation, which ultimately enables the NN to learn complex relationships between the net power injections and voltage states effectively. The application of PINN4PF is illustrated through four test power systems, namely 4-bus [39], 15-bus [40], 290-bus [41], and 2224-bus [42] test systems.
II Power Flow Analysis
Power flow (PF) analysis aims to specify the state variables of power systems, i.e., $[\delta\,v]^{T}$, where $\delta$ and $v$ denote the voltage phase angle and magnitude, respectively. The AC PF analysis problem is a representation of Kirchhoff’s laws and is formulated in rectangular coordinates [29] as:
@@ II Power Flow Analysis
This study considers cases with one reference bus and load buses for simplicity. A standard approach to handling pv buses is converting pv buses to pq buses by estimating the reactive power while applying constraints to maintain voltage levels within acceptable limits. Once converted, pv buses can be treated as pq buses for training. It should also be noted that while a voltage magnitude of $1.0$ p.u. and a phase angle of 00 degrees are assumed in this study, it is flexible to accommodate different reference bus settings.
@@
where $p_{i}$ and $q_{i}$ are defined by (1a) and (1b), respectively, and $p_{i}^{d}$ and $q_{i}^{d}$ are the net active and reactive power injections at bus $i$, respectively. (2) is conventionally solved iteratively to specify $[\delta\,v]^{T}$ until a convergence criterion is met, i.e., the mismatch between $p_{i}$ and $p_{i}^{d}$ but also $q_{i}$ and $q_{i}^{d}$ are small enough.
III Deep Learning-based Power Flow Analysis
@@
IV Proposed architecture: PINN4PF
PINN4PF includes two advancements in forward and loss steps, respectively denoted by A and B in Fig. 1, resulting in a double-head architecture enhanced with an adaptive activation function and a physics-based loss function. Note that classical PF solvers, e.g., the Newton-Raphson (NR) method, and PINN4PF can be interchangeably used to specify the state of the power system. The following subsections provide detailed information on the proposed novel advancements.
@@ IV-B Loss Function
where $N$ is the total number of data points $j$. $f(\cdot)\in\{(f_{0},f_{1},f_{2})\}$ and $f^{\prime}(\cdot)$ are the NN architecture (3a)–(3c) and the physical model (11), respectively. $\beta_{0}$ and $\beta_{1}$ represent the coefficients for the supervised and physical penalty terms, respectively, with the constraint that $\beta_{0}+\beta_{1}=1$. Initially, $\beta_{0}=1$ and $\beta_{1}=0$. After 100 epochs, the value of $\beta_{1}$ increases at a specified rate, with its maximum value determined from the sensitivity analysis. $\theta=\{W,b\}$ is the trainable parameters of the NN and $\alpha$ is the trainable parameter for the activation function.
V Results
@@
Input features and output labels of the datasets are represented as $\vec{x}$ and $\vec{y}$, respectively, where $\vec{x}$ includes $(\vec{p}^{d}_{i},\vec{q}^{d}_{i})$ and $\vec{y}$ includes $(\vec{\mu}_{i},\vec{\omega}_{i})$ obtained from the Newton-Raphson method (NR) for all load buses $i$. The PandaPower Python package [47] is used to perform NR and generate the datasets. Note that PandaPower specifies the state variables of power systems, i.e., $[\delta\text{ }v]^{T}$, that is, there is a need for converting the state variables to $\mu_{i}=v_{i}cos\delta_{i}$ and $\omega_{i}=v_{i}sin\delta_{i}$ to yield the output labels $\vec{y}=\{(\vec{\mu}_{i},\vec{\omega}_{i}):i=1,2,\dots,n\}$.
V-A Model Setup
@@ V-B1 Generalization ability
The mean and standard deviation of the mean squared error (MSE) for the test dataset obtained by PINN4PF and MLP are compared in Fig. 2. PINN4PF is observed to achieve up to 85% and 65% lower maximum testing MSE for the voltage magnitude $[V^{2}]$ and voltage phase angle $[rad^{2}]$, respectively, compared to MLP.
@@
Figure 2: Mean and standard deviation of MSE for direct physical quantities: (a) voltage magnitude $[V^{2}]$ and (b) voltage phase angle $[rad^{2}]$. The results are obtained by PINN4PF and MLP based on the testing dataset for the 15-bus test system.
V-B2 Robustness
@@
The robustness of PINN4PF, MLP, and LR is evaluated using the 15-bus test system under varying noise levels. Controlled noise levels ranging from 0% to 10% are introduced to both input features and output labels of the training dataset. The corrupted vectors are defined as $\vec{x}^{\prime}=\vec{x}\pm\vec{r}_{x}$ and $\vec{y}^{\prime}=\vec{y}\pm\vec{r}_{y}$, where $\vec{r}_{x}$ and $\vec{r}_{y}$ are vectors of random values between 0 and 1, and 0 and 0.1, respectively. Fig. 3 illustrates the changes in the MSE for the voltage magnitude $[V^{2}]$ based on the testing dataset with varying noise levels for MLP. The comparison is made relative to a constant curve, representing the MSE obtained for PINN4PF using the highest noise level, i.e., 10%. At the 0% noise level, MLP outperforms PINN4PF trained with 10% noisy data by approximately 18%. However, as the noise level increases, the performance of MLP deteriorates significantly, with MSE increasing up to six times. LR is excluded from the graph as its MSE at different noise levels is two orders of magnitude higher than that of PINN4PF. In addition, LR shows limited improvement with increasing dataset sizes and consistently underperforms compared to both PINN4PF and MLP.
@@
Figure 3: Illustration of the performance of PINN4PF and MLP based on the MSE obtained for the voltage magnitude $[V^{2}]$ under varying noise levels in the training dataset for the 15-bus test system. The MSE values are normalized relative to that of PINN4PF at the 10% noise level.
V-B3 Training dataset size
@@
The impact of the size of the training dataset on the performance of MLP is investigated using the 15-bus test system. Fig. 4 illustrates the changes in the MSE for the voltage magnitude $[V^{2}]$ based on the testing dataset with varying training dataset sizes. The comparison is made relative to a constant curve representing the MSE obtained for PINN4PF using 256 training data points. It is observed that MLP requires a training dataset twice as large as that of PINN4PF to achieve a still inferior performance. However, the performance of MLP improves with more training data, reaching a comparable level with a dataset four times larger than that of PINN4PF. This indicates that PINN4PF is more data-efficient. LR is not included in the graph as its MSE with different training dataset sizes is two orders of magnitude larger than that of PINN4PF. In addition, LR shows limited improvement with increasing training dataset sizes and consistently performs worse than PINN4PF and MLP. This indicates that the capacity of LR to capture complex relationships in PF analysis is significantly lower, and increasing the amount of training data is insufficient to bridge the performance gap.
@@
Figure 4: Illustration of the performance of PINN4PF and MLP based on the MSE obtained for the voltage magnitude $[V^{2}]$ for the 15-bus test system under different training dataset sizes. The MSE values are normalized relative to that of PINN4PF with 256 training data points.
V-B4 Accuracy of derived power flow quantities
@@
For derived physical quantities, i.e., line current, line active power, and line reactive power, the mean and standard deviation of the testing MSE obtained by PINN4PF and MLP are computed and compared in Fig. 5. It is also observed that PINN4PF achieves up to 81%, 63%, and 66% lower maximum testing MSE for the line current $[A^{2}]$, line active power $[W^{2}]$, and line reactive power $[VAR^{2}]$, respectively, compared to MLP.
@@
Figure 5: Mean and standard deviation of MSE for derived physical quantities: (a) line current $[A^{2}]$, (b) line active power $[W^{2}]$, and (c) line reactive power $[VAR^{2}]$. The results are obtained by PINN4PF and MLP based on the testing dataset for the 15-bus test system.
V-B5 Scalability

### v3

@@ Abstract
This study introduces PINN4PF, an end-to-end deep learning architecture for power flow (PF) analysis that effectively captures the nonlinear dynamics of large-scale modern power systems. The proposed neural network (NN) architecture consists of two important advancements in the training pipeline: (A) a double-head feed-forward NN that aligns with PF analysis, including an activation function that adjusts to the net active and reactive power injections patterns, and (B) a physics-based loss function that partially incorporates power system topology information through a novel hidden function. The effectiveness of the proposed architecture is illustrated through 4-bus, 15-bus, 290-bus, and 2224-bus test systems and is evaluated against two baselines: a linear regression model (LR) and a black-box NN (MLP). The comparison is based on (i) generalization ability, (ii) robustness, (iii) impact of training dataset size on generalization ability, (iv) accuracy in approximating derived PF quantities (specifically line current, line active power, and line reactive power), and (v) scalability. Results demonstrate that PINN4PF outperforms both baselines across all test systems by up to two orders of magnitude not only in terms of direct criteria, e.g., generalization ability, but also in terms of approximating derived physical quantities.
keywords:
@@ 1 Introduction
PF analysis can be performed by solving nonlinear and non-convex algebraic equations derived from the nodal balance of the net active and reactive power injections at each bus in power systems [3]. Exact analytical solutions for these equations, which also involve impedance parameters, load characteristics, and generator conditions of power systems, are impossible. Therefore, iterative numerical methods, such as the Gauss-Seidel and Newton-Raphson (NR) methods, are conventionally employed to converge to a solution that satisfies the PF equations within specified accuracy limits. Eventually, the solutions yield voltage phasors across the entire power system and, hence, provide a comprehensive understanding of the operational state of the power system [4].
@@
Several studies in the literature have investigated the impact of various modifications on the efficacy of deep learning-based PF analysis (e.g., [18, 19]). These modifications are categorized into three main stages: (i) pre-training (e.g., [20]); (ii) training (e.g., [21]); and (iii) post-training (e.g., [22]), as outlined in Table 1. A majority of these studies have primarily focused on the training stage. The table also highlights instances where topology information, such as line physical properties (e.g., [23]) and graph topology (e.g., [24]) has been integrated into the training stage. The literature review also indicates the utilization of different prior knowledge for deep learning-based PF analysis and optimal power flow (OPF), including equality and inequality constraints (e.g., [25]) and PF equations (e.g., [26]). Among the three presented stages, this work focuses on the training stage.
Table 1: List of literature on deep learning-based power flow (PF) and optimal power flow (OPF). Studies are grouped into: (i) pre-training; (ii) training; (iii) post-training.
@@
$p_{i}$=nodal active power, $q_{i}$=nodal reactive power, $s_{ij}$=line apparent power, |v||v|=voltage magnitude, $\delta$=voltage angle, $\mathcal{G}$=graph topology, $l_{ij}$=line physical properties, AA=Adjacency matrix, $p^{g}$=generator active power, $q^{g}$=generator reactive power, $p^{d}$=load active power, $q^{d}$=load reactive power.
Measurement data has been used in [20] to enhance deep learning-based OPF. While this modification has demonstrated improvements in accuracy, its applicability is confined to the pre-training stage. Physical properties have been used on a few occasions in the past to enhance deep learning-based PF analysis (e.g., [23, 36]), where the focus was mainly on the training stage. However, the high dependency on the physical properties of power systems affects the efficiency of this modification, certainly because physical properties are not always reliable due to different reasons, including aging and environmental conditions. In addition, integrating all the physical properties makes training processes computationally very expensive. The completion idea and filtering technique have also been employed in a few studies (e.g., [35]), where the focus is mainly on the post-training stage. Yet, similar to pre-training stage modifications, this approach has limitations as it is detached from the training stage. Among various methods, modified loss functions have been widely used for deep learning-based PF analysis (e.g., [27]) and OPF (e.g., [32]). Physics-based loss functions have shown significant promise in enhancing the performance and reliability of NNs for PF and OPF.
This paper proposes an end-to-end deep learning architecture, hereafter called physics-informed neural network for power flow (PINN4PF), where the focus is on the training stage and extends the existing knowledge of physics-informed NNs for PF analysis (e.g., [21, 27]). PINN4PF learns the mapping from net active ($p$) and reactive ($q$) power injections to complex bus voltages. It employs a double-head feed-forward neural network, where one head predicts the real part ($\mu$) and the other the imaginary part ($\omega$) of the complex voltage. Each head uses a modified ReLU activation function with a trainable slope, enabling adaptive learning of nonlinear relationships during training via backpropagation.
A key innovation is a physics-based loss function derived from the power balance equations, which incorporates prior knowledge of the system through a novel hidden function that partially encodes line characteristics. Unlike methods that require full system information (e.g., GNNs), our approach reduces dependency on complete topology data. Despite this, PINN4PF achieves up to two orders of magnitude improvement in performance compared to conventional NNs. The application of PINN4PF is illustrated through four test power systems, namely 4-bus [37], 15-bus [38], 290-bus [39], and 2224-bus [40] test systems. The main contributions are:
@@
    A double-head feed-forward NN for PF analysis, with adaptive activation functions responsive to the net active and reactive power injections patterns.
@@
    A physics-based loss function that partially integrates power system topology, thus reducing the need for complete line data.
2 Power Flow Analysis
Power flow (PF) analysis aims to specify the state variables of power systems, i.e., $\delta\,v]^{T}$, where $\delta$ and $v$ denote the voltage phase angle and magnitude, respectively. The AC PF analysis problem is a representation of Kirchhoff’s laws and is formulated in rectangular coordinates [27] as:
@@ 2 Power Flow Analysis
This study considers cases with one reference bus and load buses for simplicity. A standard approach to handling pv buses is converting pv buses to pq buses by estimating the reactive power while applying constraints to maintain voltage levels within acceptable limits. Once converted, pv buses can be treated as pq buses for training. It should also be noted that while a voltage magnitude of $1.0$ p.u. and a phase angle of 0 degrees are assumed in this study, it is flexible to accommodate different reference bus settings.
@@
where $p_{i}$ and $q_{i}$ are defined by (1a) and (1b), respectively, and $p_{i}^{d}$ and $q_{i}^{d}$ are the net active and reactive power injections at bus $i$, respectively. (2) is conventionally solved iteratively to specify $\delta\,v]^{T}$ until a convergence criterion is met, i.e., the mismatch between $p_{i}$ and $p_{i}^{d}$ but also $q_{i}$ and $q_{i}^{d}$ are small enough.
3 Deep Learning-based Power Flow Analysis
@@
4 Proposed architecture: PINN4PF
The proposed physics-informed neural network for power flow (PINN4PF) includes two advancements in forward and loss steps, respectively denoted by A and B in Fig. 1, resulting in a double-head architecture enhanced with an adaptive activation function and a physics-based loss function. Note that classical PF solvers, e.g., the Newton-Raphson (NR) method, and PINN4PF can be interchangeably used to specify the state of the power system. The following subsections provide detailed information on the proposed novel advancements.
@@ 4.2 Loss Function
where NN is the total number of data points $j$. $f(\cdot)\in\{(f_{0},f_{1},f_{2})\}$ and $f^{\prime}(\cdot)$ are the NN architecture (3a)–(3c) and the physical model (11), respectively. $\beta_{0}$ and $\beta_{1}$ represent the coefficients for the supervised and physical penalty terms, respectively, with the constraint that $\beta_{0}+\beta_{1}=1$. Initially, $\beta_{0}=1$ and $\beta_{1}=0$. After 100 epochs, the value of $\beta_{1}$ increases at a specified rate, with its maximum value determined from the sensitivity analysis. $\theta=\{W,b\}$ is the trainable parameters of the NN and $\alpha$ is the trainable parameter for the activation function.
5 Results
@@
Input features and output labels of the datasets are represented as $\vec{x}$ and $\vec{y}$, respectively, where $\vec{x}$ includes $(\vec{p}^{d}_{i},\vec{q}^{d}_{i})$ and $\vec{y}$ includes $(\vec{\mu}_{i},\vec{\omega}_{i})$ obtained from the Newton-Raphson method (NR) for all load buses $i$. The PandaPower Python package [45] is used to perform NR and generate the datasets. Note that PandaPower specifies the state variables of power systems, i.e., $\delta\text{ }v]^{T}$, that is, there is a need for converting the state variables to $\mu_{i}=v_{i}cos\delta_{i}$ and $\omega_{i}=v_{i}sin\delta_{i}$ to yield the output labels $\vec{y}=\{(\vec{\mu}_{i},\vec{\omega}_{i}):i=1,2,\dots,n\}$.
5.1 Model Setup
@@ 5.2.1 Generalization ability
The mean and standard deviation of the mean squared error (MSE) for the test dataset obtained by PINN4PF and MLP are compared in Fig. 2. PINN4PF is observed to achieve up to 85% and 65% lower maximum testing MSE for the voltage magnitude $V^{2}]$ and voltage phase angle $rad^{2}]$, respectively, compared to MLP.
@@
Figure 2: Mean and standard deviation of MSE for direct physical quantities: (a) voltage magnitude $V^{2}]$ and (b) voltage phase angle $rad^{2}]$. The results are obtained by PINN4PF and MLP based on the testing dataset for the 15-bus test system.
5.2.2 Robustness
@@
The robustness of PINN4PF, MLP, and LR is evaluated using the 15-bus test system under varying noise levels. Controlled Gaussian noise with standard deviations corresponding to noise levels from 0% to 10% is added to both input features and output labels of the training dataset. The corrupted vectors are defined as $\vec{x}^{\prime}=\vec{x}\pm\vec{r}_{x}$ and $\vec{y}^{\prime}=\vec{y}\pm\vec{r}_{y}$, where $\vec{r}_{x}$ and $\vec{r}_{y}$ are vectors of random values between 0 and 1, and 0 and 0.1, respectively. Fig. 3 illustrates the changes in the MSE for the voltage magnitude $V^{2}]$ based on the testing dataset with varying noise levels for MLP. The comparison is made relative to a constant curve, representing the MSE obtained for PINN4PF using the highest noise level, i.e., 10%. At the 0% noise level, MLP outperforms PINN4PF trained with 10% noisy data by approximately 18%. However, as the noise level increases, the performance of MLP deteriorates significantly, with MSE increasing up to six times. LR is excluded from the graph as its MSE at different noise levels is two orders of magnitude higher than that of PINN4PF. In addition, LR shows limited improvement with increasing dataset sizes and consistently underperforms compared to both PINN4PF and MLP.
@@
Figure 3: Illustration of the performance of PINN4PF and MLP based on the MSE obtained for the voltage magnitude $V^{2}]$ under varying noise levels in the training dataset for the 15-bus test system. The MSE values are normalized relative to that of PINN4PF at the 10% noise level.
5.2.3 Training dataset size
@@
The impact of the size of the training dataset on the performance of MLP is investigated using the 15-bus test system. Fig. 4 illustrates the changes in the MSE for the voltage magnitude $V^{2}]$ based on the testing dataset with varying training dataset sizes. The comparison is made relative to a constant curve representing the MSE obtained for PINN4PF using 256 training data points. It is observed that MLP requires a training dataset twice as large as that of PINN4PF to achieve a still inferior performance. However, the performance of MLP improves with more training data, reaching a comparable level with a dataset four times larger than that of PINN4PF. This indicates that PINN4PF is more data-efficient. LR is not included in the graph as its MSE with different training dataset sizes is two orders of magnitude larger than that of PINN4PF. In addition, LR shows limited improvement with increasing training dataset sizes and consistently performs worse than PINN4PF and MLP. This indicates that the capacity of LR to capture complex relationships in PF analysis is significantly lower, and increasing the amount of training data is insufficient to bridge the performance gap.
@@
Figure 4: Illustration of the performance of PINN4PF and MLP based on the MSE obtained for the voltage magnitude $V^{2}]$ for the 15-bus test system under different training dataset sizes. The MSE values are normalized relative to that of PINN4PF with 256 training data points.
5.2.4 Accuracy of derived power flow quantities
@@
For derived physical quantities, i.e., line current, line active power, and line reactive power, the mean and standard deviation of the testing MSE obtained by PINN4PF and MLP are computed and compared in Fig. 5. It is also observed that PINN4PF achieves up to 81%, 63%, and 66% lower maximum testing MSE for the line current $A^{2}]$, line active power $W^{2}]$, and line reactive power $VAR^{2}]$, respectively, compared to MLP.
@@
Figure 5: Mean and standard deviation of MSE for derived physical quantities: (a) line current $A^{2}]$, (b) line active power $W^{2}]$, and (c) line reactive power $VAR^{2}]$. The results are obtained by PINN4PF and MLP based on the testing dataset for the 15-bus test system.
5.2.5 Scalability